What is the pH and pOH of a solution prepared by dissolving 1.50 g of Ca(OH)2 in water to make 965 mL of solution?
Part A
What is the pH of a solution?
Express your answer using two decimal places.
pH = ???

Part B
What is the pOH of a solution?
Express your answer using two decimal places.

pOH = ???
1
Expert's answer
2015-11-17T08:53:59-0500
Solution:
Part A)
Calcium hydroxide dissociates in the solution to produce hydroxyl anions, which create the basic pH of the solution.
Ca(OH)2 = Ca2+ + 2OH-
From the given mass we can calculate the concentration of the hydroxyl groups:
n(Ca(OH)2)=m(Ca(OH)2)/M(Ca(OH)2)=1.5/74=0.02 mol
n(OH−)=2*n(Ca(OH)2)=2*0.02=0.04 mol
Because V = 965 mL = 0.965 L :
[OH−]=c(OH−)=n(OH−)/V=0.04/0.965=0.042 M
And we can estimate the pH from the calculating pOH with the help of the formulas:
pOH = - log [OH-] from this, pOH = -log (0.042) = 1.38
pOH + pH = 14 from this, pH = 14 - pOH
pH = 14-1.38 = 12.62

Part B)
[OH−] = 0.042 M
pOH = - log [OH-] from this, pOH = -log (0.042) = 1.38

Answer:
Part A) pH = 12.62
Part B) pOH = 1.38